OpenStudy (anonymous):

im missing 1 more step.help plz! 5.2a + 10)(a + 10) 2a^2 + 20a + 10a + 100 2a^2 + 30a + 100 a = - 5 and a = - 10 It’s your turn to be a game show host! As you know, in the game of Math Time, the contestants are given an answer and they must come up with the question that corresponds to the given answer. Your task for this portion of the assignment is to create two different “answers” (and the questions that accompany them) that the host could use for the final round of Math Time. The questions and answers you create must be unique.Example: Question for Host: x2 + 6x + 5 is the pro

OpenStudy (anonymous):

(x+4)(x+3)=x^2+7x+12 (x-7)(x+2)=x^2-5x-14 im i correct for part 2???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

THANK YOU=]